### v4

# LLMBind: A Unified Modality-Task Integration Framework

Abstract.
In the multi-modal domain, the dependence of various models on specific input formats leads to user confusion and hinders progress. To address this challenge, we introduce LLMBind, a novel framework designed to unify a diverse array of multi-modal tasks. By harnessing a Large Language Model (LLM) via Mixture-of-Experts (MoE), LLMBind processes multi-modal inputs and generates task-specific tokens, enabling the invocation of corresponding models to accomplish tasks. This unique approach empowers LLMBind to interpret inputs and generate outputs across various modalities, including image, text, video, and audio. Furthermore, we have constructed an interaction dataset comprising 400k instructions, which unlocks the ability of LLMBind for interactive visual generation and editing tasks. Extensive experimentation demonstrates that LLMBind achieves very superior performance across diverse tasks and outperforms existing models in user evaluations conducted in real-world scenarios. Moreover, the adaptability of LLMBind allows for seamless integration with the latest models and extension to new modality tasks, highlighting its potential to serve as a unified AI agent for modeling universal modalities. [^1] The source code will be made publicly available.
AI Agent, Multimodal Large Language Models, Mixture-of-Experts
ccs: Computing methodologies Multi-agent planningccs: Computing methodologies Computer vision tasksccs: Computing methodologies Robotic planning
1. Introduction
Figure 1. Existing Multimodal Large Language Models are currently designed in three ways: (a) to obtain understanding ability by adding an encode-side adapter; (b) to achieve generation ability by decode-side adapter or (c) by text prompt. Differently, our method achieves modality-task integration by (d) task-specific text prompts and token embedding.
Figure 2. Example Demonstrations. The LLMBind is designed to facilitate a user-friendly AI agent for achieving human-like conversation, interactive image, video, and audio generation, as well as interactive image editing and segmentation.
Recently, there has been significant progress in the field of Artificial Intelligence, including the Large Language Models (LLMs) (Chiang et al., 2023; Touvron et al., 2023) for natural language processing tasks and the diffusion models (Nichol et al., 2022; Tang et al., 2023; Liu et al., 2023a) for high-quality modality generation. Nevertheless, these models are typically limited to few modalities and tasks, highlighting the necessity to develop Multi-Modal Language Models (MLLMs) to integrate and address a broader range of tasks.
Existing methods tend to solve limited tasks and do not explore the unification of more tasks within a shared LLM. As shown in Fig. 1, we first use two characteristics to evaluate the efficacy of a framework, the updateability which quantifies the ease of replacing the frozen task-specific model, and the expandability which indicates the ability of the framework to extend to other tasks. Then, we summarize the existing MLLMs into three categories: 1) Encode-side Adapter, such as MiniGPT-4 (Zhu et al., 2023a) and LLaVA (Liu et al., 2023c), typically utilize encode-side adapters to handle inputs beyond language and can only perform the visual question-answering task; 2) Decode-side Adapter (Wu et al., 2023; Ghosal et al., 2023; Koh et al., 2023), as shown in Fig. 1(b), incorporates decode-side adapters for various generation models, enhancing the adaptation of LLMs to complex multimodal tasks like generation. However, these methods lack updateability as they require training a new adapter when introducing a new generation model; 3) Text Prompt Agent (Xia et al., 2023; Brooks et al., 2023), as depicted in Fig. 1(c), can directly generate desired text prompts for specific modality tasks without additional training. However, these approaches can only perform tasks that only have text prompts as additional inputs, and cannot solve more complex visual tasks such as image segmentation.
In this work, to achieve both updateability and expandability, we introduce LLMBind, a unified framework to integrate models for various modality tasks. As depicted in Fig. 1(d), LLMBind is capable of processing two modalities: image and text. It processes these inputs through a visual encoder (Radford et al., 2021) and LLM (Chiang et al., 2023), ultimately producing task-specific text prompts or token embeddings. For tasks such as video generation (Ding et al., 2021b; Rombach et al., 2022a), image generation (Nichol et al., 2022; Tang et al., 2023), audio generation (Liu et al., 2023a; Yang et al., 2023b), and image editing (Brooks et al., 2023; Fu et al., 2023), we utilize text prompts as input to existing pre-trained generation models, significantly reducing the training complexity of our framework and enhancing its updateability. For modality tasks like image segmentation (Lai et al., 2023; Liu et al., 2023b), classification (Yang et al., 2023a; Naeem et al., 2023), and detection (Ge et al., 2021) that cannot be solely addressed with text prompts, we extract relevant semantic visual information from images using learnable task-specific token embeddings. Subsequently, these embeddings are utilized as an auxiliary input to the corresponding modality-task models.
To guide the LLM to generate task-specific text prompts and token embeddings, we expand the vocabulary of LLM with learnable task-specific tokens, which can be divided into two categories: Task-Prompt Tokens and Semantic-Embedding Tokens. By training the LLM with specially designed instruction data pairs and task-specific tokens, we can instruct the model to discriminate the task type from user input and generate the corresponding text prompt or token embedding for a given multimodal task. Moreover, to better unify different modality tasks within the shared LLM, we introduce a highly effective technique called Mixture-of-Experts (MoE), which has been used successfully in recent studies (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Riquelme et al., 2021; Ahmed et al., 2016; Fedus et al., 2022). By incorporating this technique, our LLMBind can handle various modality tasks by utilizing different experts to collaboratively process multi-modal tasks, as shown in Fig. 2.
In practical applications, users often require multi-turn interactive generation and editing to achieve precise and satisfactory results. For example, in image editing, users need to apply various effects repeatedly to achieve their desired visual outcome. To this end, we develop a dataset comprising 400k instructional samples for interactive generation and editing tasks. Each sample, generated by ChatGPT, is formatted as a multi-round dialogue. This dataset alleviates the scarcity of instruction data for generation and editing tasks, equipping LLMBind with robust interaction capabilities.
Extensive experiments have demonstrated LLMBind’s superior performance in tasks such as referring segmentation, video generation, image generation, and audio generation, surpassing numerous specialized and multitask models. Additionally, LLMBind has exhibited remarkable performance in human evaluations of interactive generation tasks, exceeding traditional generation methods in terms of both the quality of generated data and its alignment with text. Overall, LLMBind demonstrates promising results in advancing the development of human-like MLLM and AI agents. Our Contributions to this work are as follows:
  • We introduce LLMBind, a unified framework for various modality tasks, encompassing image segmentation, generation, editing, as well as video and audio generation. We employ the MoE technique to effectively handle diverse tasks.
  • We devise a task-specific token strategy to seamlessly integrate different tasks into a unified MLLM, ensuring high model’s updateability and expandability for emerging tasks.
  • We curate an interactive dataset comprising 400k instructional data, enabling LLMBind to tackle multi-turn interactive visual generation and editing tasks.
2. Related Work
2.1. Cross-Modal Generation and Editing
Multi-modal generation has remained a highly active research area in recent years. These models (Ramesh et al., 2022; Rombach et al., 2022b; Hong et al., 2022a; Kong et al., 2020; Yang et al., 2023c) allow the generation of corresponding images, videos, and audio based on specified text prompts. However, this approach requires users to directly write text prompts for image generation, lacking interactive capabilities. Another approach to cross-modal generation (Wu et al., 2023; Zeqiang et al., 2023), combines dialogue with generation, making it more user-friendly and aligned with current user needs. Given that generated images may not always meet user requirements, image editing has emerged as another active research area within the field of artificial intelligence-generated content. Representative works in this area include Text2live (Bar-Tal et al., 2022), Imagic (Kawar et al., 2023), Sdedit (Meng et al., 2021).
2.2. Cross-Modal Segmentation
With the introduction of frameworks such as LanguageBind (Zhu et al., 2023b) and ImageBind (Girdhar et al., 2023), which combine multi-modal data in a shared semantic space, several works focusing on cross-modal understanding (Liu et al., 2023c; Zhu et al., 2023a; Yin et al., 2023; Jin et al., 2023) have emerged. Along with these advancements, there has been a growing interest in text-guided image referring segmentation, a new approach that enables users to interact with images using user instruction. Unlike traditional semantic segmentation, which assigns class labels to pixels, referring segmentation (Lai et al., 2023; Liu et al., 2023b) aims to segment target objects based on complex textual instructions.
Table 1. Comparison of functionalities among different MLLMs. indicates capability, indicates limitation, and “-” hints at the potential for easy extension. We cover a range of tasks including text conversation (Text-Con), image comprehension (Img-Com), image generation (Img-Gen), image editing (Img-Edit), video generation (Vid-Gen), audio generation (Aud-Gen), image segmentation (Img-Seg), and image detection (Img-Det).
MLLMs Text-Con Img-Com Img-Gen Img-Edit Vid-Gen Aud-Gen Img-Seg Img-Det LLaMA (Touvron et al., 2023) Vicuna (Chiang et al., 2023) Flaminggo (Alayrac et al., 2022) BLIP2 (Li et al., 2023a) mPLUG-Owl (Ye et al., 2023) DetGPT (Pi et al., 2023) LLaVA (Liu et al., 2023c) Mini-GPT4 (Zhu et al., 2023a) Audiogpt (Huang et al., 2023) Otter (Li et al., 2023b) LISA (Lai et al., 2023) GLaMM (Rasheed et al., 2023) NExTGPT (Wu et al., 2023) LLMBind (Ours) -
2.3. Mixture-of-Experts
The concept of Mixture-of-Experts (MoE) in artificial neural networks has evolved. Early methods activated all experts, leading to high computational requirements (Eigen et al., 2013). Subsequent research introduced sparse MoE models, where distinct experts are selectively activated using a learnable gating mechanism (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022). These approaches have shown success in computer vision (Eigen et al., 2013; Ahmed et al., 2016; Gross et al., 2017; Wang et al., 2020) and natural language processing (Shazeer et al., 2017; Lepikhin et al., 2020; Zhou et al., 2022) domains.
2.4. Multimodal Large Language Models
Recently, the emergence of Large Language Models (LLM) has ushered AI into a new era. With the open-sourcing of LLM models like LLaMA (Touvron et al., 2023) and Vicuna (Chiang et al., 2023), numerous research works have successfully explored methods to extend LLMs to other modalities, including video, audio, and images. This has led to the development of a variety of related MLLM models, as shown in Table 1. For instance, Flamingo (Alayrac et al., 2022) employs cross-attention structures for visual-text interaction, while DetGPT (Pi et al., 2023) connects multimodal LLMs with detectors, enabling object detection based on user instructions. BLIP-2 (Li et al., 2023a) encodes input images into queries using the Q-Former structure. mPLUG-Owl (Ye et al., 2023) and LLaVA (Liu et al., 2023c) encode images into image features using image encoders, which are then combined with text embeddings for LLM learning. LISA (Lai et al., 2023) achieves referring segmentation by training LLMs to output mask token embeddings in an end-to-end manner. NExTGPT (Wu et al., 2023) implements multimodal generation by aligning the output of the adapter, which is added after the LLM, with the output text features from the text encoder of various generation models (Liu et al., 2023a; Rombach et al., 2022a).
3. Methodology
In this work, we present LLMBind, a unified framework for model integration, aiming to handle various modality tasks into a shared LLM. In Section 3.1, we detail the components of LLMBind and the process of their integration. In Section 3.2, we discuss our implementation of the LoRA Mixture-of-Experts technique. In Section 3.3, we focus on the training objectives and the specific loss functions we use for model optimization. Finally, we describe the dataset crucial for the training of LLMBind in Section 3.4.
3.1. LLMBind
Overview. The framework of LLMBind is illustrated in Fig. 3. It primarily consists of a vision transformer (Radford et al., 2021) as the image encoder, a large language model called Vicuna (Chiang et al., 2023), a SAM-like mask decoder as well as image, audio, and video diffusion models (Rombach et al., 2022a; Tang et al., 2023; Singer et al., 2022b) for generation and editing tasks.
Figure 3. LLMBind architecture. Utilizing task-specific tokens, our framework implements universal multimodal understanding, segmentation, editing, and generation capabilities. During training, we employ LoRA-MoE techniques to facilitate model learning of different modality tasks and reduce computational complexity.
Visual Encoder and Tokenizer. We prioritize the processing of inputs from two primary modalities: images and text. For image encoding, we employ the ViT-L/14 (Radford et al., 2021) to extract image features. We first reshape the image $x\in\mathbb{R}^{H\times W\times C}$ into a sequence of flattened 2D patches $X_{v}=[x_{v}^{1},x_{v}^{2},...,x_{v}^{N}]\in\mathbb{R}^{N\times(P^{2}\times C)}$, where $(H,W)$ is the shape of the original image, $C$ is the number of channels, $N=\frac{H\times W}{P^{2}}$ is the resulting number of image patches. The LLM uses constant latent vector size $D$ through all of its layers, so we use two MLP layers to map the image patches to $D$ dimensions. Similarly, text is processed through a text embedding function $g$, which projects it into a sequence of text tokens $X_{t}=[x_{t}^{1},x_{t}^{2},\ldots,x_{t}^{N}]\in\mathbb{R}^{N\times D}$, where $N$ is the length of this text token sequence.
Task-Specific Tokens. Different multimodal tasks often require distinct network structures, making it difficult to consolidate them within a shared LLM. To overcome this challenge, we expand the vocabulary of the LLM with specially designed, learnable task-specific tokens. As shown in Fig. 4 , these tokens can be classified into two categories: Task-Prompt Tokens and Semantic-Embedding Tokens. The first category includes tokens similar to standard text tokens, such as <gen> and </gen> for image generation, <vid_cap> and </vid_cap> for video generation, and <edit> and </edit> for image editing. These tokens appear in pairs, and the content between them represents the intermediate text prompt required for the corresponding modality task. In contrast, the second category is imbued with dense semantic information, diverging from traditional text tokens. For example, <seg> contains rich visual information related to segmented objects in an image. After decoding by a mask decoder, this token can be converted back into masks. This category also encompasses tokens like <detect> and <cls>, designed for extracting spatial coordinates and classification data in image-based tasks. This dual-token framework not only enhances the adaptability of the LLM across various modalities but also ensures scalability and extensibility.
Figure 4. Task-Specific tokens include two categories: Task-Prompt Tokens and Semantic-Embedding Tokens.
Modality Task Integration. We seamlessly integrate various modality-task models in our framework using task-specific text prompts and token embeddings generated by LLM. For understanding tasks, we directly output text. In the generation and editing scenarios, we feed text prompts into the corresponding task model. Specifically, we use Stable Diffusion (Rombach et al., 2022a) for image generation, CoDi (Tang et al., 2023) for audio generation, ModelScopeT2V (Wang et al., 2023) for video generation, Instruct-pix2pix (Brooks et al., 2023) for image editing, thus ensuring high-quality, modality-specific output. For complex tasks such as referring segmentation, we use a single task-specific token embedding to learn the semantic information of the image features with respect to the segmentation instructions, which is then processed by a dedicated segmentation decoder. This Integration method is adaptable to various task requirements and has good extensibility.
3.2. LoRA Mixture-of-Experts
To enable effective learning for different multimodal tasks through collaboration among diverse experts, we introduce a LoRA Mixture-of-Experts (MoE) technique. In the traditional transformer architecture, for a given input $x$, the forward propagation process of each decoder block can be written as follows:
(1) $f(x)=x+f_{FFN}(x),$
where $f_{FFN}$ denotes the feed-forward network. The matrix operation of the linear layer in the FFN block can be expressed as:
(2) $o=Wx=W_{0}x+\triangle Wx,$
where $W_{0}\in\mathbb{R}^{d_{in},d_{out}}$ denotes the parameter metrix, $\triangle W\in\mathbb{R}^{d_{in},d_{out}}$ represents the parameter update and $o$ is the output of the linear layer. After we substitute the linear layer with the MoE technique, the forward process of the MoE layer can be expressed as follows:
(3) $o=W_{0}x+\triangle Wx=W_{0}x+\sum_{i=1}^{N}G(x)_{i}E_{i}(x),$
(4) $G(x)=Softmax(x\cdot W_{g}),$
where $E_{i}(\cdot)$ denotes the $i$-th expert, $G(\cdot)$ represents the router network in the MoE layer and $W_{g}$ is a trainable parameter. Using the router network, these parallel experts can efficiently handle different multimodal tasks through collaboration. To save training cost, following LoRA-MoE (Dou et al., 2023), we adapt the experts in the MoE layer into a low-rank format and the matrix $\triangle W_{E}\in\mathbb{R}^{d_{in}\times d_{out}}$ of a single expert can be expressed as follows:
(5) $\triangle W_{E}=BA,$
where $A\in\mathbb{R}^{d_{in}\times r}$, $B\in r\times\mathbb{R}^{d_{in}}$ and the rank $r<<min(d_{in},d_{out})$. The forward process of the second modified linear layer can be written as follows:
(6) $o=W_{0}x+\gamma\sum_{i=1}^{N}G(x)_{i}E_{i}(x)=W_{0}x+\frac{\alpha}{r}\sum_{i=1}^{N}{w_{i}}B_{i}A_{i}x,$
where $w_{i}$ is the attention weight of $i$-th expert and $\alpha$ is a constant hyper-parameter. This low-rank format Mixture-of-Experts greatly reduces the computational overhead imposed by MoE, resulting in a reduction of training parameters to approximately 0.3% of the total parameter count.
3.3. Training Objective
The training of our model involves three objectives: text auto-regressive loss $\mathcal{L}_{reg}$, segmentation mask loss $\mathcal{L}_{mask}$ and MoE auxiliary loss $\mathcal{L}_{aux}$. These objectives are combined using weighted coefficients ($\lambda_{reg}$, $\lambda_{mask}$ and $\lambda_{aux}$) to compute overall objective $\mathcal{L}$:
(7) $\mathcal{L}=\lambda_{reg}\mathcal{L}_{reg}+\lambda_{mask}\mathcal{L}_{mask}+\lambda_{aux}\mathcal{L}_{aux}.$
Auto-Regressive Loss. For task-specific text prompt training, we optimize the output of LLM through a generative loss in an auto-regressive manner. Given input tokens $X$, we compute the auto-regressive loss by:
(8) $\mathcal{L}_{\text{reg}}(\theta)=\left\{\begin{array}[]{lr}-\sum_{i=1}^{N}log(p_{\theta}(x_{i}|X_{v},X_{t,<i})),&\text{if}\ X_{v}\subseteq X\\ \\ -\sum_{i=1}^{N}log(p_{\theta}(x_{i}|X_{t,<i})),&\text{if}\ X_{v}\nsubseteq X\\ \end{array}\right.$
where $\theta$ is the trainable parameter, $X_{v}$ are the image tokens, $X_{t,<i}$ are the instruction and answer tokens in all turns before the current prediction token $x_{i}$.
Segmentation Loss. To compute the segmentation loss $\mathcal{L}_{mask}$, a combination of per-pixel Binary Cross-Entropy loss $\mathcal{L}_{BCE}$ and Die Coefficient loss $\mathcal{L}_{DICE}$ is used, with corresponding loss weights, $\lambda_{bce}$ and $\lambda_{dice}$. Given the ground-truth targets $M$ , the loss can be expressed as follows:
(9) $\mathcal{L}_{mask}=\lambda_{bce}\mathcal{L}_{BCE}(\hat{{M}},{M})+\lambda_{dice}\mathcal{L}_{DICE}(\hat{{M}},{M}),$
where $\hat{{M}}$ represents the predicted segmentation masks.
MoE Auxiliary Loss. To ensure a balanced load across experts, we integrate an auxiliary loss  (Fedus et al., 2022). For each MoE layer, this auxiliary loss is added to the total model loss during training. With $N$ experts numbered from $i=1$ to $N$, and considering a batch $\mathcal{B}$ consisting of $T$ tokens, we determine this auxiliary loss through the scaled dot-product of two vectors, $f$ and $P$, as follows:
(10) $\mathcal{L}_{aux}=\alpha\cdot N\cdot\sum_{i=1}^{N}f_{i}\cdot P_{i},$
where $\alpha$ is a multiplicative coefficient, $f_{i}$ represents the proportion of tokens assigned to expert $i$, calculated by:
(11) $f_{i}=\frac{1}{T}\sum_{x\in\mathcal{B}}\mathbbm{1}\{\text{argmax}\>p(x)=i\}.$
Meanwhile, $P_{i}$ indicates the fraction of the router probability allocated for expert $i$.
(12) $P_{i}=\frac{1}{T}\sum_{x\in\mathcal{B}}p_{i}(x).$
Here, $p_{i}(x)$ is the probability of routing token $x$ to expert $i$.
Table 2. Referring segmentation results on three referring segmentation datasets: refCOCO (Kazemzadeh et al., 2014), refCOCO+ (Kazemzadeh et al., 2014), and refCOCOg (Mao et al., 2016) with metric cIoU. The best results and second best results are indicated by boldface and underline, respectively.
Method refCOCO refCOCO+ refCOCOg
val testA testB val testA testB val(U) test(U)
MCN (Luo et al., 2020) 62.4 64.2 59.7 50.6 55.0 44.7 49.2 49.4
VLT (Ding et al., 2021a) 67.5 70.5 65.2 56.3 61.0 50.1 55.0 57.7
CRIS (Wang et al., 2022) 70.5 73.2 66.1 62.3 68.1 53.7 59.9 60.4
LAVT (Yang et al., 2022) 72.7 75.8 68.8 62.1 68.4 55.1 61.2 62.1
ReLA (Liu et al., 2023b) 73.8 76.5 70.2 66.0 71.0 57.7 65.0 66.0
X-Decoder (Zou et al., 2023a) - - - - - - 64.6 -
SEEM (Zou et al., 2023b) - - - - - - 65.7 -
LISA (Lai et al., 2023) 74.9 79.1 72.3 65.1 70.8 58.1 67.9 70.6
LLMBind (ours) 76.9 78.5 73.2 67.8 71.9 60.8 69.8 70.8
Table 3. Text-to-image generation results on COCO-captions dataset (Lin et al., 2014), evaluated with the Fréchet Inception Distance (FID) for image quality.
Method FID ($\downarrow$)
Specialized Models
CogVideo (Ding et al., 2021b) 27.10
GLIDE (Nichol et al., 2022) 12.24
CoDi (Tang et al., 2023) 11.26
Multitask Models
NExT-GPT (Wu et al., 2023) 11.28
LLMBind-SD1.5 16.41
LLMBind-SD2 14.29
LLMBind-GALIP* 10.38
Figure 5. Different modality task datasets used in LLMBind, including VQA, Interactive Generation (video, audio, and image), Interactive Image Editing, and Segmentation.
Figure 6. Prompts for ChatGPT to construct interactive image generation and editing data.
3.4. Dataset
The training datasets consist of both publicly available datasets and our custom proprietary ones, as visualized in  Fig. 5. These proprietary datasets are specifically constructed through interactive processes enabled by ChatGPT.
Interactive Image Generation and Editing Dataset. To empower the model in executing image generation and editing tasks, following Mini-DALLE3 (Zeqiang et al., 2023), we design a prompt for ChatGPT to construct 400k human-like interactive instruction data, tackling the problem of image generation and editing through textual description. As shown in  Fig. 6, we first clarify the role and capability of ChatGPT within the prompt. We provide it with instruction text formats, tags, and examples for both generating and editing tasks. However, we notice that the generated samples closely resemble the examples in terms of text format, tone, and content. To address this, we further clarify the diversity of the target data in terms of the above multiple aspects in the prompt. Following a manual review of a portion of our dataset, we have confirmed that this approach is highly effective in ensuring significant diversity.
Interactive Video and Audio Generation Dataset. To construct video and audio generation instruction data, we adapt the t2x(Wu et al., 2023) dataset and change its format to resemble that used for the image generation task. Some examples are listed in Appendix C.
Segmentation Datasets. For segmentation task, we incorporate several semantic segmentation datasets, including ADE20K (Zhou et al., 2017), COCO-Stuff (Caesar et al., 2018), PACO-LVIS (Ramanathan et al., 2023), PartImageNet (He et al., 2022), and PASCAL-Part (Chen et al., 2014). We also utilize referring segmentation datasets like refCLEF, refCOCO, refCOCO+(Kazemzadeh et al., 2014), and refCOCOg(Mao et al., 2016).
Table 4. Text-to-audio generation results on AudioCaps dataset (Kim et al., 2019), evaluated with FD for audio similarity and IS for quality and diversity.
Method FD ($\downarrow$) IS ($\uparrow$)
Specialized Models
DiffSound (Yang et al., 2023b) 47.68 4.01
AudioLDM-S (Liu et al., 2023a) 29.48 6.90
AudioLDM-L (Liu et al., 2023a) 23.31 8.13
Multitask Models
NExT-GPT (Wu et al., 2023) 23.58 8.35
LLMBind-CoDi 22.90 8.77
LLMBind-Auffusion 21.99 10.57
(a) Text to video generation
(b) Image generation and edition
(c) Image segmentation
(d) Text to audio generation
Figure 7. Demonstration examples of the proposed LLMBind framework on different modality tasks.
Visual Question Answering Datasets. We incorporates the LLaVA-Instruct-665k dataset (Liu et al., 2023d) for visual question-answering task. This dataset enables the training of models to comprehend, reason, and make decisions accurately based on visual content.
4. Experiments
4.1. Experimental Setting
Training Specifications. During training, we employ 8 NVIDIA A100 GPUs (40G). The training scripts are based on the deepspeed (Rasley et al., 2020) engine. We use the AdamW (Loshchilov and Hutter, 2017) optimizer with a learning rate of 0.0003 and weight decay of 0. The WarmupDecayLR learning rate scheduler is adopted, with a warmup iteration count of 100. The auto-regressive loss and segmentation loss have weights of 1.0 each. The BCE loss and DICE loss have weights of 2.0 and 0.5 respectively. We set the batch size per device to 2, and accumulate gradients over 10 steps, with a total of 15k steps.
Data Details. During training, to sustain the stability of multi-task training, we configure the data ratios for Segmentation data, Visual Question Answering data, and Interactive Generation and Editing instruction data at a 1:1:1 ratio. Within the segmentation dataset, the proportions for semantic segmentation, referring segmentation, and reasoning segmentation are established at 9:3:1. In the interactive instructions, the distribution for tasks like image generation, image editing, audio generation, and video generation is balanced at a 1:1:1:1 ratio. We train jointly on all task datasets to ensure that our model avoids loss fluctuations and balances losses across tasks.
Evaluation Metrics. Our experiments utilize some key metrics for precise evaluation. Text-to-audio generation is assessed using Fréchet Distance (FD) for distribution similarity and Inception Score (IS) for quality and diversity. Text-to-video generation leverages Fréchet Inception Distance (FID) for content quality and CLIPSIM for textual alignment. In text-to-image generation, FID evaluates image distribution similarity. For segmentation, gIoU is defined by the average of all per-image Intersection-over-Unions (IoUs), while cIoU is defined by the cumulative intersection over the cumulative union. In addition, to better showcase the model’s generative capabilities during user interaction, we employed human evaluation to assess the quality of the modality data generated and its alignment with the given text instructions.
Table 5. Text-to-image generation results on COCO-captions dataset (Lin et al., 2014), evaluated with the Fréchet Inception Distance (FID) for image quality.
Method FID ($\downarrow$)
Specialized Models
CogVideo (Ding et al., 2021b) 27.10
GLIDE (Nichol et al., 2022) 12.24
CoDi (Tang et al., 2023) 11.26
Multitask Models
NExT-GPT (Wu et al., 2023) 11.28
LLMBind-SD1.5 16.41
LLMBind-SD2 14.29
LLMBind-GALIP* 10.38
4.2. Results Visualization
As shown in  Fig. 7, LLMBind can be executed on various modal tasks such as video, audio, image generation, image editing, and image segmentation, providing users with a more user-friendly interactive experience.
4.3. Quantitative Analysis
Referring Segmentation. For referring segmentation task, as illustrated in Table 2, LLMBind achieves superior results on various datasets. On the refCOCO dataset, for the val, testA, and testB splits, LLMBind achieves scores of 76.9, 78.5, and 73.2, respectively. Similarly, on the refCOCO dataset, the model records scores of 67.8, 71.9, and 60.8 for val, testA, and testB. On the refCOCOG dataset, LLMBind achieves 69.8 and 70.8.
Image Generation. In the text-to-image generation task, as shown in Table 5, LLMBind utilizes various baseline models, achieving impressive FID scores on the COCO-caption (Lin et al., 2014) dataset. Specifically, with Stable Diffusion v1.5 and v2.0 (Rombach et al., 2022c) as baselines, LLMBind attains FID scores of 16.41 and 14.29, respectively. Notably, using GALIP (Tao et al., 2023) as the baseline model, LLMBind secures an FID score of 10.38, outperforming models such as NeXT-GPT (Wu et al., 2023) and CoDi (Tang et al., 2023).
Audio Generation. In the domain of text-to-audio generation, LLMBind has been meticulously evaluated and compared with existing models. As depicted in  Table 4, LLMBind utilizes the CoDi (Tang et al., 2023) model as a baseline for audio generation, attaining an FD score of 22.90 and an IS of 8.77 on the AudioCaps (Kim et al., 2019) dataset. Furthermore, when employing Auffusion (Tang et al., 2023) as the foundational model for generation, LLMBind secures an FD score of 21.99 and an IS score of 10.57 on the same dataset, demonstrating a discernible advantage over models such as NeXT-GPT (Wu et al., 2023) and AudioLDM (Liu et al., 2023a).
Table 6. Text-to-video generation results on MSR-VTT dataset (Xu et al., 2016), evaluated with FID for video quality and CLIPSIM for video coherence.
Method FID ($\downarrow$) CLIPSIM ($\uparrow$)
Specialized Models
CogVideo (Hong et al., 2022b) 23.59 0.2631
Latent-VDM (Rombach et al., 2022a) 14.25 0.2756
Latent-Shift (An et al., 2023) 15.23 0.2773
Make-A-Video (Singer et al., 2022a) 13.17 0.3049
Multitask Models
NExT-GPT (Wu et al., 2023) 13.04 0.3085
LLMBind-Show-1 13.08 0.3072
LLMBind-ModelScopeT2V 11.09 0.2930
LLMBind-HiGen 8.60 0.2947
Table 7. Human evaluation statistics regarding the preference rate for Quality Score (Q.S.) and Alignment Score (A.S.). The former metric assesses the quality of generated content, while the latter evaluates the alignment between the generated data and textual instructions. Both metrics are rated on a scale of 1-4, with the final score being the weighted average.
Method Q.S. Final A.S. Final
1 2 3 4 score 1 2 3 4 Score
Audio Generation
Direct input 7% 35% 34% 24% 2.75 18% 39% 23% 20% 2.45
Ours 8% 28% 35% 29% 2.85 6% 24% 27% 43% 3.07
Image Generation
Direct input 5% 39% 43% 13% 2.64 21% 29% 45% 5% 2.34
Ours 5% 27% 41% 27% 2.90 12% 19% 51% 18% 2.75
Video Generation
Direct input 29% 41% 23% 7% 2.08 23% 42% 21% 14% 2.26
Ours 24% 38% 27% 11% 2.25 19% 40% 24% 17% 2.39
Video Generation. For the text-to-video generation task, LLMBind’s performance on the MSR-VTT (Xu et al., 2016) dataset is shown in Table 6. Using ModelScopeT2V (Wang et al., 2023) as the base model, LLMBind achieves an FID of 11.09 and a CLIPSIM score of 0.2930. With Show-1 (Zhang et al., 2023) as the base model, it obtains an FID of 13.08 and a CLIPSIM score of 0.3072. When employing HiGen (Qing et al., 2023) as the base model, LLMBind reaches an FID of 8.60 and a CLIPSIM score of 0.2947, showing comparable performance to NExT-GPT (Wu et al., 2023). As LLMBind avoids lossy training on generation models, its generative power primarily relies on the selected base models.
4.4. Human Evaluation
As shown in  Fig. 8, to assess the practical performance of LLMBind in interactive Image, Video, and Audio generation tasks (IVA), we conducted a comprehensive user study. This study is designed to measure the capability of models to generate IVA data in response to interactive text instructions, emulating natural human-like communication. The study focuses on assessing two critical metrics: Quality Score (Q.S.) and Alignment Score (A.S.). The Q.S. assesses the quality of the generated IVA content, while A.S. evaluates the alignment between the generated data and the textual instructions provided. Participants interact with the model by issuing text instructions and then rate each generated sample on a scale of 1 to 4 for both metrics, with 1 representing poor performance and 4 indicating excellent performance. The aggregated results from a panel of 30 participants, as illustrated in  Table 7, clearly indicate that LLMBind excels in both the quality of modal data generation and its ability to align with text-based instructions closely. The final scores, calculated as the weighted average of the ratings for Q.S. and A.S., demonstrate LLMBind’s proficiency in understanding and executing complex, multimodal tasks as directed by human users.
Figure 8. Human evaluation. The evaluation indicators for interactive generation include Q.S. and A.S.
Table 8. Ablation study about the rank of LoRA.
LoRA Training MoE Expert Top-K gIoU cIoU
RANK Steps Layers Nums Experts
4 7k 16/32 4 2 59.4 62.5
8 7k 16/32 4 2 58.9 65.8
12 7k 16/32 4 2 58.7 66.2
16 7k 16/32 4 2 58.6 65.2
Table 9. Ablation study about the training steps.
LoRA Training MoE Expert Top-K gIoU cIoU
RANK Steps Layers Nums Experts
8 5k 16/32 4 2 58.7 59.4
8 6k 16/32 4 2 58.6 62.2
8 7k 16/32 4 2 58.9 65.8
8 8k 16/32 4 2 59.4 65.8
Table 10. Ablation study about the number of MoE layers.
LoRA Training MoE Expert Top-K gIoU cIoU
RANK Steps Layers Nums Experts
8 7k 16/32 4 2 59.9 65.8
8 7k 10/32 4 2 59.2 60.9
8 7k 8/32 4 2 58.9 57.1
8 7k 6/32 4 2 58.3 61.8
Table 11. Ablation study about the number of experts and the number of Top-K activated experts.
LoRA Training MoE Expert Top-K gIoU cIoU
RANK Steps Layers Nums Experts
8 7k 16/32 4 2 59.4 66.5
8 7k 16/32 3 2 58.9 65.8
8 7k 16/32 2 2 58.7 66.2
8 7k 16/32 2 1 58.6 65.2
4.5. Ablation Analysis
We apply the ablation analysis on the reasoning segmentation task, since it is conducive to quantitative analysis and comparison. The detailed tables are listed in the supplementary.
Rank of LoRA. As shown in  Table 8, our framework exhibits slight performance variation across different settings.
Training Steps. As shown in  Table 9, we reveal a clear trend that longer iteration steps consistently yield improvements in both gIoU and cIoU metrics.
Number of MoE Layers. As shown in  Table 10, model performance generally improves with the increase in the number of MoE layers.
Number of All Experts and Top-K Activated Experts. As shown in  Table 11, when the number of experts is fixed, more activated experts improve model performance. Conversely, increasing the number of experts enhances performance when the number of activated experts is constant.
5. Conclusion and Future Direction
In this work, we introduce the LLMBind framework, designed to address the limitations of current Multimodal Large Language Models in handling a wide range of multi-modality tasks. LLMBind efficiently integrates tasks across various modalities such as image segmentation, video generation, and more, through its unique Task-Specific tokens. To enhance the adaptability and efficiency of our framework, we introduce a novel dataset of 400k interactive instructions and an innovative Mixture-of-Experts architecture. Extensive experimental validation demonstrates that LLMBind excels in a variety of tasks and holds immense potential for expansion to other modality tasks.
Our framework marks a significant step towards creating a unified AI agent capable of handling universal modalities. Looking ahead, we aim to further expand the LLMBind framework to include more modality tasks while concurrently focusing on investigating and refining training strategies for multi-task joint training.
 Appendix for LLMBind
Appendix A Interactive Architecture
A.1. Multi-Turn and Interactive
LLMBind allows users to accomplish various modality-tasks as mentioned above through a dynamic and interactive conversational approach. Unlike the traditional fixed-format prompt inputs, users have more freedom to express their requirements and can improve the model’s outputs through interactive dialogue, such as modifying the content of generated images. Additionally, LLMBind can automatically preprocess the user’s input text intent and analyze the corresponding modality-task, greatly reducing the user’s learning curve and enhancing the overall user experience.
A.2. Interactive Multi-modal Generation
LLMBind can automatically determine the corresponding data type for generation based on the details provided in the input text. This includes image generation, video generation, and audio generation. Unlike traditional generation methods that require specific text descriptions, LLMBind leverages the capabilities of LLM to create image captions that align with user-provided objects, scenes, or concepts. This assists users in conceptualizing image content and achieving their desired image generation outcome. For image generation, LLMBind generates textual descriptions that match the user’s intent by incorporating <gen>DETAILED IMAGE CAPTION</gen> in the output response. Subsequently, LLMBind employs a generation model prompt style using “DETAILED IMAGE CAPTION” to generate the corresponding image. It is worth noting that during this process, users are not required to adhere to a fixed prompt format. For video and text generation, LLMBind follows a similar approach, replacing <gen>,</gen> with <vid_cap>,</vid_cap> and <aud_cap>,</aud_cap> respectively.
A.3. Interactive Image Editing and Refinement
LLMBind offers interactive image editing and refinement capabilities. Image editing typically involves modifying image content based on textual instructions. This can involve significant changes to objects and attributes within the image. The edited image may differ greatly from the original, but it closely aligns with the user’s textual instructions. On the other hand, image refinement entails making adjustments and optimizations to enhance the details of the image. The edited image should maintain a significant resemblance to the original but incorporates specific image adjustments based on the provided text instructions.
A.4. Interactive Text Conversation
Text-based conversation is a potential capability of LLMs, and LLMBind excels in maintaining effective communication while simultaneously supporting other modality tasks. This functionality enables seamless switching between different modality tasks and highlights the user-friendly nature of the model.
Appendix B Impact Statements
B.1. Broader Impacts
This study explores the development of our LLMBind model in the field of artificial intelligence, particularly its multimodal capabilities. The development and implementation of the model could have significant impacts on multiple areas of society, both positive and potentially negative.
  • Creative Industries and Intellectual Property: The LLMBind model’s capabilities in content generation have far-reaching implications for the creative industries. It can facilitate the democratisation of content creation and the innovation process, but at the same time may challenge existing norms of intellectual property and originality, raising concerns about authors’ rights and creative ethics.
  • Information Integrity and the Media: The model’s ability to edit and generate realistic content could have implications for the journalism and media sectors. While being able to automate and improve the accuracy of information dissemination, there is also a risk of being used to create misleading or false information, posing a threat to the authenticity of public discourse.
  • Human-AI Interaction and Social Implications: As the LLMBind model advances in dialogue, its integration in everyday life may affect social dynamics and interpersonal interactions. While they can provide companionship and support in therapeutic and educational settings, they also raise questions about privacy, emotional attachment, and reduced interpersonal interactions.
B.2. Reproducibility
We have ensured the availability of all models and source codes in a public domain. The provided materials include code that enables the replication of our results.
B.3. Licenses
The majority of this project is released under the Apache 2.0 license.
Appendix C Demo and Training Datasets
Figure 9. Example of Image Generation and Editing (1).
Figure 10. Example of Image Generation and Editing (2).
Figure 11. Example of Image Generation and Editing (3).
Figure 12. Example of Image Generation and Editing (4).
Figure 13. Example of Image Segmentation (1).
Figure 14. Example of Image Segmentation (2).
Figure 15. Example of Image Segmentation (3).
Figure 16. Examples of Video Generation.
Figure 17. Examples of Audio Generation.
From Value
HUMAN I’ve been thinking about getting a tattoo of a cosmic jellyfish, can you help me visualize it?
GPT Of course! Here’s an image of a cosmic jellyfish: <gen> a cosmic jellyfish with neon tentacles floating in space </gen>
HUMAN Wow, that’s amazing! Can you show me a cosmic jellyfish with some galaxies in the background?
GPT Certainly! Here’s a modified image of the cosmic jellyfish with galaxies in the background: <edit> a cosmic jellyfish with neon tentacles floating amidst colorful galaxies in space </edit>
HUMAN This is exactly what I had in mind! I’m so excited about this tattoo now. Can you show me a smaller version of the cosmic jellyfish, maybe in a minimalist style?
GPT Absolutely! Here’s a minimalist version of the cosmic jellyfish: <gen> a simple line art of a cosmic jellyfish with a serene and modern aesthetic </gen>
HUMAN I love it! It’s going to be hard to choose which one to get. Thanks for all your help!
GPT You’re welcome! I’m glad I could assist you in visualizing your cosmic jellyfish tattoo. Let me know if there’s anything else I can do for you.
HUMAN I’m curious about a magical forest with glowing mushrooms. Can you show me what it looks like?
GPT Certainly! Here is the concept of a magical forest with glowing mushrooms: <gen> a magical forest with glowing mushrooms </gen>
HUMAN Wow, that looks amazing! Is there a special creature that lives in this forest?
GPT Yes, there is! Let me show you a unique creature that inhabits this magical forest: <edit> a magical forest with glowing mushrooms and a special creature </edit>
HUMAN I’d love to see more details of the creature and its surroundings. Can you provide a closer look?
GPT Absolutely! Here is a closer look at the special creature and its enchanting surroundings: <edit> a magical forest with glowing mushrooms, featuring a special creature in its habitat </edit>
HUMAN This is fantastic! Can you tell me more about the magical properties of the glowing mushrooms?
GPT Certainly! The glowing mushrooms in this forest possess magical qualities, emitting a soft and enchanting light that illuminates the entire area, creating an otherworldly atmosphere.
HUMAN I’m planning a trip to the beach; can you show me what a picturesque beach looks like?
GPT Sure! Here is a breathtaking beach scene. <gen> a picturesque beach with golden sand and crystal-clear blue water </gen>
HUMAN Wow, that looks amazing! Can you also show me a beach with palm trees and hammocks?
GPT Absolutely! Here is a serene beach setting with palm trees and hammocks. <edit> a picturesque beach with palm trees and hammocks </edit>
HUMAN This looks so relaxing. I wonder what a beach at sunset looks like. Can you show me that?
GPT Of course! Here is a stunning sunset beach view. <edit> a picturesque beach at sunset </edit>
HUMAN Oh, that’s so beautiful. I’d love to have this as a desktop wallpaper. Can you create one for me?
GPT Sure thing! Here is a desktop wallpaper featuring a picturesque beach at sunset. <edit> a desktop wallpaper of a picturesque beach at sunset </edit>
Table 12. Interactive image generation and editing data
From Value
HUMAN Can you find me an audio recording that showcases the progression of a motor’s sound? I’m particularly interested in capturing the transition from a humming idle to a revving acceleration.
GPT Certainly! I have just the audio clip you’re looking for. It begins with the calm hum of the motor as it idles, and then gradually builds up to a roaring acceleration. This recording beautifully captures the evolving soundscape of the motor’s performance. <audio_cap> A motor hums as it idles then accelerates </audio_cap>
HUMAN I’m in need of an audio file that captures the thrilling experience of an engine revving, with the added excitement of humming and squealing tires. Can you help me with that?
GPT Absolutely! I’ve got the perfect audio snippet for you. It’s a recording that captures the adrenaline-pumping sound of an engine revving, accompanied by the intense humming and the screeching of tires. Get ready to feel the rush! <audio_cap> Revving of an engine with humming and squealing tires </audio_cap>
HUMAN Can you create an audio clip with the sound of people speaking while a car engine runs?
GPT Absolutely! I have just the audio clip you’re looking for. It captures the ambient noise of a lively conversation happening alongside the constant hum of a running car engine. The combination of human voices and the mechanical sound of the engine creates an intriguing auditory experience. Enjoy! <audio_cap> People speaking while car engine runs </audio_cap>
HUMAN I’d love to listen to a recording of a duck quacking loudly with insect noises in the background.
GPT Absolutely! I have just the audio clip for you. Prepare to be immersed in nature’s orchestra as you hear the distinct sound of a duck quacking loudly, accompanied by the gentle symphony of buzzing insects in the background. Enjoy the harmonious blend of these natural sounds! <audio_cap> A duck quacking loudly with insect noises in the background </audio_cap>
HUMAN I need some audio that captures the essence of a thunderstorm, with the added element of a vehicle in motion. Can you help me find it?
GPT Of course! Prepare to be transported into the heart of a thunderous storm. I’ve got the perfect audio recording for you: the symphony of rain cascading from the sky, intermingled with the rolls of thunder, all while a vehicle races through the rain-soaked streets. Get ready to feel the power and energy of nature in this captivating audio experience. <audio_cap> Rain falls and thunder rolls while a vehicle runs </audio_cap>
HUMAN I’m trying to recreate a specific audio ambiance in my story. It involves a man speaking, and in the background, you can hear birds chirping intermittently. Additionally, there’s a small machine that makes short whirring sounds. Can you help me paint this audio picture?
GPT Absolutely! Picture a tranquil scene outdoors, where the delicate melodies of chirping birds dance in the air. Amidst this natural symphony, a man’s voice resonates, creating an engaging narrative. Accompanying his speech are brief intervals of a small machine emitting gentle whirring sounds, adding a touch of mechanical presence to the serene environment. It’s a unique audio blend that brings together the beauty of nature and human activity. <audio_cap> A man speaks and a small machine whirs shortly several times while birds chirp in the background </audio_cap>
HUMAN I’m in the mood for some soothing sounds. Can you find me an audio recording of water flowing in a stream?
GPT Of course! I have just the audio you’re looking for. Here’s a recording that captures the tranquil melody of water gently flowing in a serene stream. Enjoy the peaceful ambiance! <audio_cap> Water is flowing in a stream </audio_cap>
Table 13. Interactive audio generation data
From Value
HUMAN I’m curious to see a video where someone switches a light on and off repeatedly. Can you find one for me?
GPT Absolutely! I found a video that captures the act of toggling a light switch on and off repeatedly. It should satisfy your curiosity and provide an interesting visual representation of this action. <video_cap> Turning light on and off </video_cap>
HUMAN I’m curious to see a video where someone switches a light on and off repeatedly. Can you find one for me?
GPT Absolutely! I found a video that captures the act of toggling a light switch on and off repeatedly. It should satisfy your curiosity and provide an interesting visual representation of this action. <video_cap> Turning light on and off </video_cap>
HUMAN I have a dusty floor that needs cleaning. I wonder what it would look like if a vacuum cleaner cleaned the floor during the daytime.
GPT Ah, I can certainly help you visualize that. Imagine a bright and sunny day with rays of sunlight streaming through the windows. The vacuum cleaner glides across the floor, sucking up the dust particles, leaving behind beautifully clean and clear surfaces. The sound of the vacuum hums in the background, indicating the effectiveness of its cleaning. Here’s a video clip capturing the scene for you to get a better sense of it. <video_cap> Vacuum cleaner cleaning the floor with dust on it during daytime </video_cap>
HUMAN Show me a video that encapsulates the allure of Venice during its carnival season. I want to witness the joyful interactions between tourists and the carnival masks in Piazza San Marco.
GPT Your wish is my command! Here’s a captivating video that perfectly captures the magic of Venice’s carnival. It showcases the enchanting atmosphere of Piazza San Marco, where tourists mingle with the carnival masks, creating moments of pure joy and excitement. <video_cap> Venice - carnival masks are photographed with tourists in piazza san marco </video_cap>
HUMAN I could really use some uplifting music right now. Can you find a video of a happy man in headphones, listening to cheerful music and dancing in an armchair in a cozy room at home?
GPT Absolutely! Here’s a video that perfectly captures the essence of joy and relaxation. Watch as a young man in casual wear, sitting in an armchair in a cozy room at night, enjoys his weekend by listening to cheerful music on his smartphone. His infectious happiness will surely brighten up your day. <video_cap> Happy man in headphones listens to cheerful music on smartphone sitting and dancing in armchair in cozy room at home at night. young male person in casual wear is relaxing on weekend. 4k static shot </video_cap>
HUMAN I’d like to see a video of cute adorable children having fun together. It would be great if they were isolated on a white background.
GPT Absolutely! I have just the video in mind. Get ready to witness the joyous sight of cute and adorable children engaging in fun activities, all set against a pristine white background. Enjoy the heartwarming moments of pure happiness and laughter! <video_cap> Cute adorable children having fun together isolated on white background </video_cap>
HUMAN I’m in need of some visual tranquility. Is there a video that showcases a waterfall cascading over moss-covered rocks, with the camera gradually panning towards a lush green landscape?
GPT Absolutely! I have the perfect video to soothe your soul. Watch as water gently trickles on mossy rocks beneath a captivating waterfall, creating a serene atmosphere. The camera slowly pans to the right, unveiling a breathtaking view of a lush and vibrant green landscape that will transport you to a state of pure tranquility. <video_cap> Water trickles on mossy rocks underneath a waterfall, panning right towards a lush green landscape </video_cap>
Table 14. Interactive video generation data